OpenStudy (anonymous):

14ab3 ---------- 7a^-2b^-1 Which of the following is the simplest form of the expression above?Select one of the options below as your answer: A. 2a-2b-3 B. 2a-1b2 C. 2a-1/2b-3 D. 2a3b

OpenStudy (jackellyn):

2x^3b^4

ganeshie8 (ganeshie8):

@jackellyn how you got that... can you pls show steps ?

OpenStudy (jackellyn):

14/7 =2 When you divide exponents you're supposed to subtract them For the a exponent a-(-2a) =3 For the b exponent 3- (-1)=4 Therefore, 2a^3b^4

ganeshie8 (ganeshie8):

hey got that.. i was working with wrong expression lol. .thanks

OpenStudy (jackellyn):

No problem :)